# Linear complementary dual quasi-cyclic codes of index 2

### v2

@@ Abstract
Keywords: Linear codes, Euclidean LCD codes, Hermitian LCD codes, symplectic LCD codes, quasi-cyclic codes
@@ 2.2 Decomposition of quasi-cyclic codes
For each $i$ and $j$, let $F_{i}=F[x]/(f_{i})$, $H_{j}^{\prime}=F[x]/(h_{j})$, and $H_{j}^{\prime\prime}=F[x]/(h_{j}^{*})$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity over $F$. Let $\xi^{u_{i}}$ and $\xi^{v_{j}}$ be roots of $f_{i}(x)$ and $h_{j}(x)$, respectively. Then we also have $h_{j}^{*}(\xi^{-v_{j}})=0$, and $F_{i}\cong F(\xi^{u_{i}})$, $H_{j}^{\prime}\cong F(\xi^{v_{j}})$, and $H_{j}^{\prime\prime}\cong F(\xi^{-v_{j}})$, see [18, p. 136].
@@
where $\textbf{c}=(c_{1}(\xi^{u_{i}}),\dots,c_{\ell}(\xi^{u_{i}})),\textbf{d}=(d_{1}(\xi^{u_{i}}),\dots,d_{\ell}(\xi^{u_{i}}))\in F_{i}^{\ell}.$ This is the inner product induced by $x\mapsto x^{-1}$, not the usual Hermitian inner product, see [28, p. 2693]. For each $1\leq j\leq p$, the vector space $(H_{j}^{\prime})^{\ell}\cong(H_{j}^{\prime\prime})^{\ell}$ is equipped with the usual Euclidean inner product and $\perp_{e}$} denotes the usual Euclidean dual.
@@
Since $f_{i}(x)$ is self-reciprocal, the cardinality of $F_{i}$, say $q_{i}$, is an even power of $q$ for all $1\leq i\leq s$ with two exceptions. One of these exceptions, for all $m$ and $q$, is the field coming from the irreducible factor x1x-1 of $x^{m}-1.$ When $q$ is odd and $m$ is even, $x+1$ is another self-reciprocal irreducible factor of $x^{m}-1$. In these cases, $q_{i}=q$. Except for two cases, the Hermitian inner product $\langle\cdot,\cdot\rangle_{h}$ is equivalent to the usual Hermitian inner product, see also [17, p.72] and [18, p.136]. For the two exceptions, in which case the corresponding field $F_{i}$ is $F$, we equip $F_{i}^{\ell}$ with the usual Euclidean inner product. Then the previous formula for $\langle\cdot,\cdot\rangle_{h}$ is still true, since $\xi^{u_{i}}=\xi^{-u_{i}}=\pm 1$.
We have the following characterization of Euclidean LCD QC codes from [18, Theorem 7.3.6].
@@ 2.3 Quasi-cyclic codes of index 2
In [36, Lemma 1], Séguin showed that if $C$ is a quasi-cyclic code generated by one element $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid x^{m}-1$ and $g(x)=\gcd(g_{11}(x),g_{12}(x))$, then $\dim C=m-\deg g(x)$. With the choice of $g_{22}(x)$ described in Lemma 2.4, we see that the dimension of $C$ in Theorem 2.2 is consistent with the result by Séguin.
@@ 1. Since $g$ and $l$ are self-reciprocal,
In [8], the authors showed that any linear code is equivalent to a Euclidean LCD code for $q>3$. Using Theorem 3.1, we construct several binary ($q=2$) (in Table 1) and ternary $(q=3)$ quasi-cyclic Euclidean LCD codes.
@@
$[46,23,10]$ $x+1$ $1+x^{2}+x^{6}+x^{8}$ $x^{9}+x^{13}+x^{15} +$$x^{20}$ $(x^{11}+x^{10}+x^{6}+x^{5}$ $x^{4}+x^{2}+1)$$(x^{11}$ $x^{9}+x^{7}+x^{6} +$$x^{5}+x+1)$ Near-optimal
$[50,24,10]$ $(x+1)(x^{4}+x^{3}$ $x^{2}+x+1)$ $(x+1)(1+x^{5}+$ $x^{7}+x^{14})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1) -$
$[50,28,8]$ $x+1$ $(x+1)(1+x^{5}$ $x^{7}+x^{8})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1)$ BKLC
$[54,21,12]$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{2}+x+1)$ $(x^{4}+x+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ -
$[54,25,11]$ $(x+1)$ $(x^{2}+x+1)$ $(x^{2}+x+1)$ $(x^{13}+x^{9}+x^{7}$ $x^{6}+x^{5}+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ New
$[54,26,8]$ $(x+1)$ $(x^{2}+x+1)$ $(x+1)$ $(x^{8}+x^{7}+x^{5}+1)$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ New
Table 1: Binary quasi-cyclic (Euclidean) LCD codes, where BKLC means that the minimum distance of the code is equal to the minimum distance of Best Known Linear Code according to the Grassl’s table, “New” means codes are new to the codes given in [14], [24], [44] with given parameters. “Near-optimal” means that the minimum distance is one less than the optimal distance for linear codes (it might be optimal as an LCD code), and it is better than the LCD code [46,23,9] in [44].
Example 1.
Let $q=3$ and $m=13$. Then $x^{13}-1=(x+2)(x^{3}+2x+2)(x^{3}+x^{2}+2)(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2)$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x^{3}+2x+2,\\ g_{12}(x)&=x^{7}+2x^{4}+x+2,\\ g_{22}(x)&=(x^{3}+x^{2}+2)(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2).\end{split}$
Then, by Theorem 3.1, $C$ is a quasi-cyclic ternary LCD code with parameters $[26,14,7]$. Its minimal distance is the same as that of the Best Known Linear Code (BKLC) as per Grassl’s Table. The dual code has parameters $[26,12,8]$.
Example 2.
Let $q=3$ and $m=13$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{5}+2x^{4}+2x^{3}+x^{2}+x+1,\\ g_{22}(x)&=(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2).\end{split}$
Then, by Theorem 3.1, $C$ is a quasi-cyclic ternary LCD code with parameters $[26,19,4]$. This is a near optimal code as per Grassl’s Table.
@@ 2. The conditions $\gcd(g_{11}^{\prime},g_{12})=1$ and $\gcd(g_{11}^{\prime},\bar{g}_{12})=1$ imply that
This result is consistent with that in [4] and [14].

### v3

@@ Abstract
Keywords: Linear codes, Euclidean LCD codes, Hermitian LCD codes, symplectic LCD codes, quasi-cyclic codes.
Mathematics subject classification: 94B05, 94B15, 94B60.
@@ 1 Introduction
All computations in this paper have been done with the computer algebra system MAGMA [6].
@@ 2.2 Decomposition of quasi-cyclic codes
For each $i$ and $j$, let $F_{i}=F[x]/(f_{i})$, $H_{j}^{\prime}=F[x]/(h_{j})$, and $H_{j}^{\prime\prime}=F[x]/(h_{j}^{*})$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity over $F$. Let $\xi^{u_{i}}$ and $\xi^{v_{j}}$ be roots of $f_{i}(x)$ and $h_{j}(x)$, respectively. Then we also have $h_{j}^{*}(\xi^{-v_{j}})=0$, and $F_{i}\cong F(\xi^{u_{i}})$, $H_{j}^{\prime}\cong F(\xi^{v_{j}})$, and $H_{j}^{\prime\prime}\cong F(\xi^{-v_{j}})$, see [24, p. 136].
@@
where $\textbf{c}=(c_{1}(\xi^{u_{i}}),\dots,c_{\ell}(\xi^{u_{i}})),\textbf{d}=(d_{1}(\xi^{u_{i}}),\dots,d_{\ell}(\xi^{u_{i}}))\in F_{i}^{\ell}.$ This is the inner product induced by $x\mapsto x^{-1}$, not the usual Hermitian inner product, see [33, p. 2693]. For each $1\leq j\leq p$, the vector space $(H_{j}^{\prime})^{\ell}\cong(H_{j}^{\prime\prime})^{\ell}$ is equipped with the usual Euclidean inner product and $\perp_{e}$} denotes the usual Euclidean dual.
@@
Since $f_{i}(x)$ is self-reciprocal, the cardinality of $F_{i}$, say $q_{i}$, is an even power of $q$ for all $1\leq i\leq s$ with two exceptions. One of these exceptions, for all $m$ and $q$, is the field coming from the irreducible factor x1x-1 of $x^{m}-1.$ When $q$ is odd and $m$ is even, $x+1$ is another self-reciprocal irreducible factor of $x^{m}-1$. In these cases, $q_{i}=q$. Except for two cases, the Hermitian inner product $\langle\cdot,\cdot\rangle_{h}$ is equivalent to the usual Hermitian inner product, see also [23, p.72] and [24, p.136]. For the two exceptions, in which case the corresponding field $F_{i}$ is $F$, we equip $F_{i}^{\ell}$ with the usual Euclidean inner product. Then the previous formula for $\langle\cdot,\cdot\rangle_{h}$ is still true, since $\xi^{u_{i}}=\xi^{-u_{i}}=\pm 1$.
We have the following characterization of Euclidean LCD QC codes from [24, Theorem 7.3.6].
@@ 2.3 Quasi-cyclic codes of index 2
In [40, Lemma 1], Séguin showed that if $C$ is a quasi-cyclic code generated by one element $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid x^{m}-1$ and $g(x)=\gcd(g_{11}(x),g_{12}(x))$, then $\dim C=m-\deg g(x)$. With the choice of $g_{22}(x)$ described in Lemma 2.4, we see that the dimension of $C$ in Theorem 2.2 is consistent with the result by Séguin.
@@ 1. Since $g$ and $l$ are self-reciprocal,
We now describe some examples we obtained from Theorem 3.1. Following [48], we denote
$d(n,k)_{\mathbb{F}_{q}}=\max\{d(C):C\text{ is an }[n,k,d(C)]\text{ linear code over }\mathbb{F}_{q}\},$
and
$d_{\mathrm{LCD}}(n,k)_{\mathbb{F}_{q}}=\max\{d(C):C\text{ is an }[n,k,d(C)]\text{ LCD code over }\mathbb{F}_{q}\}.$
We call a linear code with parameters $n,k,d(n,k)_{\mathbb{F}_{q}}]$ an optimal code. A comprehensive record of $d(n,k)_{\mathbb{F}_{q}}$ can be found in the online database codetables.de, see  [17]. An LCD code with parameters $n,k,d_{\mathrm{LCD}}(n,k)_{\mathbb{F}_{q}}]$ is called an optimal LCD code.
In [11], Carlet et al. showed that any linear code is equivalent to a Euclidean LCD code for $q>3$. This motivates us to search for binary and ternary Euclidean LCD codes with good parameters. Using Theorem 3.1, we construct several optimal binary Euclidean LCD codes, as presented in Tables 1 and 2.
We use the abbreviation BKLC to indicate that a code is the best known linear code according to [17], although it may not be optimal. Similarly, BKLCD denotes the best known LCD code according to [29] and [48], while it also may not be optimal.
The codes obtained in Table 1 are compared with the recently updated tables of bounds on the minimum distance of binary Euclidean LCD codes in [29] and [48]. In particular, we refer to [29, Tables 4 and 5] for codes of length $29\leq n\leq 40$, and to [48, Table 6] for length $41\leq n\leq 50$.
To the best of our knowledge, there is no existing database of LCD codes with length $n>50$ in the literature. Consequently, the codes obtained in Table 2 are compared with the linear code database available at codetables.de, see [17]. We note that the BKLC codes from [17] used for comparison are not LCD codes.
@@
$30,15,7]_{2}$ $x^{2}+x+1$ $x^{12}+x^{10}+x^{9}+x$ $(x+1)(x^{4}+x+1)$ $(x^{4}+x^{3}+1)(x^{4}+x^{3}+x^{2}+x+1)$ Optimal LCD, see [29]
$30,16,6]_{2}$ $x+1$ $(x+1)$ $(x^{10}+x^{5}+x^{3}+x^{2})$ $(x+1)(x^{4}+x+1)$ $(x^{4}+x^{3}+1)(x^{4}+x^{3}+x^{2}+x+1)$ Optimal LCD, see [29]
$34,17,8]_{2}$ $x+1$ $x^{15}+x^{14}+x^{13}+x^{9}+x^{6}$ $x^{5}+x^{4}+x^{3}+x^{2}+1$ $(x^{8}+x^{5}+x^{4}+x^{3}+1)$ $(x^{8}+x^{7}+x^{6}+x^{4}+x^{2}+x+1)$ Optimal LCD, see [29]
$34,25,4]_{2}$ $x+1$ $x^{7}+x^{6}+x^{5}+x^{4}$ $x^{8}+x^{5}+x^{4}+x^{3}+1$ Optimal LCD, see [29]
$34,26,4]_{2}$ 11 $x^{6}+x^{5}+x^{3}+x^{2}$ $x^{8}+x^{5}+x^{4}+x^{3}+1$ Optimal LCD, see [29]
$42,32,4]_{2}$ $x^{2}+x+1$ $(x+1)(x^{2}+x+1)$ $(x^{4}+x^{3}+1)$ $(x^{2}+x+1)$ $(x^{3}+x+1)(x^{3}+x^{2}+1)$ Optimal LCD, see [48]
$46,23,10]_{2}$ $x+1$ $1+x^{2}+x^{6}+x^{8}$ $x^{9}+x^{13}+x^{15}+x^{20}$ $(x^{11}+x^{9}+x^{7}+x^{6}+x^{5}+x+1)$ $(x^{11}+x^{10}+x^{6}+x^{5}+x^{4}+x^{2}+1)$ BKLCD, improves [48]
$50,24,10]_{2}$ $(x+1)(x^{4}+x^{3}$ $x^{2}+x+1)$ $(x+1)(1+x^{5}+$ $x^{7}+x^{14})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1)$ $d=d_{\text{BKLCD}}$, see [48]
$50,25,10]_{2}$ $x+1$ $x^{22}+x^{21}+x^{20}+x^{18}+x^{14}$ $x^{13}+x^{12}+x^{4}+x^{3}+x$ $(x^{4}+x^{3}+x^{2}+x+1)$ $(x^{20}+x^{15}+x^{10}+x^{5}+1)$ $d=d_{\text{BKLCD}}$, see [48]
$50,28,8]_{2}$ $x+1$ $(x+1)(1+x^{5}$ $x^{7}+x^{8})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1)$ $d=d_{\text{BKLCD}}$, see [48]
Table 1: Binary quasi-cyclic Euclidean LCD codes from Theorem 3.1, length $30\leq n\leq 50$.
$C$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Comparison with [17]
$54,22,12]_{2}$ $(x+1)$ $(x^{6}+x^{3}+1)$ $x(x+1)^{3}(x^{6}+x^{3}+1)$ $(x^{10}+x^{9}+x^{6}+x^{3}+x^{2}+x+1)$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-1$
$54,25,11]_{2}$ $(x+1)$ $(x^{2}+x+1)$ $(x^{2}+x+1)$ $(x^{13}+x^{9}+x^{7}+x^{6}+x^{5}+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-1$
$54,26,11]_{2}$ $x^{2}+x+1$ $x^{3}(x+1)(x^{2}+x+1)$ $(x^{3}+x+1)(x^{6}+x^{5}+x^{4}+x+1)$ $(x^{7}+x^{6}+x^{3}+x+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-1$
$54,27,11]_{2}$ $x^{2}+x+1$ $x^{22}+x^{21}+x^{15}+x^{11}+x^{10}$ $x^{8}+x^{7}+x^{3}+x^{2}+1$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}$
$54,28,9]_{2}$ 11 $x^{23}+x^{19}+x^{12}+x^{10}+x^{9}$ $x^{8}+x^{7}+x^{5}+x^{2}+x$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-1$
$54,29,9]_{2}$ 11 $x^{23}+x^{21}+x^{19}+x^{18}+x^{16}$ $x^{10}+x^{6}+x^{4}+x^{3}+x^{2}$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-1$
$54,33,6]_{2}$ $x^{2}+x+1$ $x^{17}+x^{12}+x^{11}+x^{3}$ $(x+1)$ $(x^{18}+x^{9}+1)$ $d=d_{\text{BKLC}}-2$
$62,30,10]_{2}$ $(x+1)$ $(x^{5}+x^{2}+1)$ $(x^{5}+x^{3}+1)$ $(x+1)^{2}$ $(x^{15}+x^{14}+x^{13}+x^{9}+x^{8}+x^{7}$ $x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1)$ $(x+1)$ $(x^{5}+x^{3}+x^{2}+x+1)$ $(x^{5}+x^{4}+x^{2}+x+1)$ $(x^{5}+x^{4}+x^{3}+x+1)$ $(x^{5}+x^{4}+x^{3}+x^{2}+1)$ $d=d_{\text{BKLC}}-2$
$62,40,8]_{2}$ $x+1$ $x^{3}(x+1)^{2}$ $(x^{14}+x^{13}+x^{12}+x^{11}$ $x^{10}+x^{9}+x^{8}+x^{7} +$$x^{4}+x^{3}+x^{2}+x+1)$ $(x+1)$ $(x^{5}+x^{3}+x^{2}+x+1)$ $(x^{5}+x^{4}+x^{2}+x+1)$ $(x^{5}+x^{4}+x^{3}+x+1)$ $(x^{5}+x^{4}+x^{3}+x^{2}+1)$ $d=d_{\text{BKLC}}$
$66,30,12]_{2}$ $(x+1)$ $(x^{2}+x+1)$ $(x^{10}+x^{7}+x^{5}+x^{3}+1)$ $x^{4}(x+1)^{6}$ $(x^{2}+x+1)$ $(x^{10}+x^{7}+x^{5}+x^{3}+1)$ $(x+1)$ $(x^{2}+x+1)$ $(x^{10}+x^{7}+x^{5}+x^{3}+1)$ $(x^{10}+x^{9}+x^{8}+x^{7}+x^{6}+x^{5}$ $x^{4}+x^{3}+x^{2}+x+1)$ $d=d_{\text{BKLC}}-3$
Table 2: Binary quasi-cyclic Euclidean LCD codes from Theorem 3.1, length $54\leq n\leq 66.$
We also describe some examples of ternary quasi-cyclic Euclidean LCD codes in Table 3. We note that the codes with length $n=26$ and $n=28$ from [17] used for comparison are not LCD codes.
$C$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Remark
$14,7,6]_{3}$ $x+2$ $2x^{5}+2x^{4}+x^{3}+2$ $x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1$ Optimal, see [17]
$16,8,6]_{3}$ $x+1$ $2x^{6}+x^{5}+2x^{4}+2x^{2}$ $(x+2)(x^{2}+1)(x^{2}+x+2)(x^{2}+2x+2)$ Optimal, see [17]
$22,11,7]_{3}$ $x+2$ $x^{9}+2x^{7}+2x^{6}+x^{4}+2x^{2}$ $(x^{5}+2x^{3}+x^{2}+2x+2)$ $(x^{5}+x^{4}+2x^{3}+x^{2}+2)$ $d=d_{\text{BKLCD}}$, see [29]
$26,13,7]_{3}$ $(x+2)(x^{3}+2x+2)$ $2x^{8}+x^{6}+2x^{3}+2x^{2}$ $(x^{3}+x^{2}+2)$ $(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2)$ $d=d_{\text{BKLC}}-1$, see [17]
$26,14,7]_{3}$ $x^{3}+2x+2$ $x^{7}+2x^{4}+x+2$ $(x^{3}+x^{2}+2)(x^{3}+x^{2}+x+2)$ $(x^{3}+2x^{2}+2x+2)$ $d=d_{\text{BKLC}}$, see [17]
$26,19,4]_{3}$ $x+2$ $x^{5}+2x^{4}+2x^{3}+x^{2}+x+1$ $(x^{3}+x^{2}+x+2)$ $(x^{3}+2x^{2}+2x+2)$ $d=d_{\text{BKLC}}-1$, see [17]
$28,14,8]_{3}$ $x+2$ $x^{3}(x+2)^{4}$ $(x^{3}+2x+1)$ $(x+2)(x^{6}+x^{5}+x^{4}+x^{3}+x^{2}+x+1)$ $(x^{6}+2x^{5}+x^{4}+2x^{3}+x^{2}+2x+1)$ $d=d_{\text{BKLC}}-1$, see [17]
Table 3: Ternary quasi-cyclic Euclidean LCD codes from Theorem 3.1, length $14~\leq~n~\leq~28.$
@@ 2. The conditions $\gcd(g_{11}^{\prime},g_{12})=1$ and $\gcd(g_{11}^{\prime},\bar{g}_{12})=1$ imply that
This result is consistent with that in [5] and [19]. We also describe some examples of binary and ternary one-generator quasi-cyclic Euclidean LCD codes in Table 4.
$C$ $g_{11}(x)$ $g_{12}(x)$ Remark
$30,13,8]_{2}$ $x^{2}+x+1$ $x^{14}+x^{10}+x^{7}+x^{6}+x^{5}+x^{4}+x+1$ Optimal LCD, see [29]
$30,14,8]_{2}$ $x+1$ $x^{14}+x^{13}+x^{12}+x^{6}+x^{5}+x^{3}+x+1$ Optimal LCD, see [29]
$34,16,8]_{2}$ $x+1$ $x^{16}+x^{14}+x^{13}+x^{10}+x^{9}+x^{5}+x^{2}+x$ $d=d_{\text{BKLC}}$, see [17]
$42,18,10]_{2}$ $x^{3}+1$ $x^{19}+x^{18}+x^{17}+x^{16}+x^{12}$ $x^{11}+x^{10}+x^{9}+x^{3}+x$ $d=d_{\text{BKLCD}}$, see [48]
$42,19,10]_{2}$ $x^{2}+x+1$ $x^{20}+x^{19}+x^{18}+x^{17}+x^{16}+x^{12}$ $x^{11}+x^{8}+x^{7}+x^{5}+x^{3}+x^{2}+x+1$ $d=d_{\text{BKLCD}}$, see [48]
$42,20,10]_{2}$ $x+1$ $x^{19}+x^{18}+x^{17}+x^{13}+x^{12}$ $x^{9}+x^{8}+x^{7}+x+1$ $d=d_{\text{BKLCD}}$, see [48]
$16,7,6]_{3}$ $x+1$ $2x^{6}+x^{5}+x^{3}+2x^{2}+2x$ Optimal, see [17]
$20,9,7]_{3}$ $x+1$ $2x^{8}+x^{7}+2x^{6}+2x^{5}+x^{3}$ $d=d_{\text{BKLCD}}$, see [29]
$22,10,8]_{3}$ $x+2$ $x^{10}+2x^{8}+x^{7}+2x^{6}+x^{5}$ + $2x^{4}+2x^{2}+x$ Optimal LCD, see [29]
Table 4: Binary and ternary one-generator quasi-cyclic Euclidean LCD codes from Theorem 4.3.